A balloon that has a static charge on its surface hangs on a string. A positively charged rod is brought near to the balloon and it feels an attractive pulling force of 13 N in the positive xdirection. A slight breeze pushes the balloon with a force of 21 N at 125° in standard position. Assuming the mass of the string and balloon are negligible, what is the net force on the balloon?


1
Expert's answer
2021-01-05T12:02:56-0500
"F=\sqrt{13^2+21^2+2(13)(21)\cos{125}}\\F=17.2\ N"